Tucker Electronic System's current balance sheet shows total common equity of $3,125,000. The company has 125,000 shares of stock outstanding, and they sell at a price of $52.50 per share. By how much do the firm's market and book values per share differ? a. $30.32 b. $27.50 c. $28.88 d. $31.83 e. $33.43
Answer #1

Calculation of book value per share

Book value per share= value of equity/ number of shares

Book value per share= 3125000/125000= 25

Market value per share= 52.50

Difference= market value per share- book value per share

Difference= 52.50-25= 27.50

So correct answer is b)$27.50